the rate of reaction triples when temperature changes from 20celcious to 50celcious. Energy of activation for the reaction is 

Grade:8

1 Answers

Arun
6 years ago

The Arrhenius equation is log10k2k1=Ea(R×2.303)[(T2T1)(T1T2)]

Given:  k2k1 = 3; R=8.314 JK-1 mol–1; T1 = 20 + 273 = 293 K

and   T2 = 50 + 273 = 323 K

Subtracting the given values in the Arrhenius equation,

log10 3 = Ea(8.314×2.303)[(323  293)(323 × 293)]

Ea = (2.303 × 8.314×323 × 293×0.477)/30

= 28811.8 J mol-1

= 28.8118 kJ mol.-1